# CycleGAN-Driven Transfer Learning for Electronics Response Emulation in High-Purity Germanium Detectors

### v1

@@ 3.3 Post processing and Network Training
Normalizing and aligning the pulses is crucial for training the network. The simulated pulses are already normalized between 00 and $1$. The raw data pulses are normalized by dividing by the $80\%$ of the average of the last five samples. Pulses are then vertically shifted so that the average of the first $200$ samples is zero. This ensures that all pulses have their preamplifier decay tails and their baselines aligned, which can help the model learn the features better. We used the $99.9\%$ rise time to horizontally align the pulses. We found this method more effective for training the network than aligning by the start of the rising edge, as the pulse shape at the end of the rise time is far more consistent between events than the shape of the beginning of the charge drift, which depends on the position of the energy deposition. The simulated pulses are padded to ensure that there are $400$ samples on both sides of the $99.9\%$ rise time. The preamplifier decay time constant $\tau$ of the data pulses is calculated by the slope of a linear fit of the logarithm of the last $300$ pulse samples. Events with poor fit quality $\chi^{2}$ or anomalous values of $\tau$ are used to identify pile-up pulses in the data and remove them from training.
@@
The training and validation of CPU-Net are conducted in PyTorch [28]. Suppose we start with a simulated pulse $X$ and data pulse $X^{\prime}$. Let $\Lambda(X)$ be the output of the REN generator and $\bar{\Lambda}(X^{\prime})$ be the output of the IREN generator, both for any pulse $Z$. During training, a simulated pulse $X$ is first fed to $\Lambda$ to produce a translated pulse $\Lambda(X)$. The discriminator $\delta_{T}$ attempts to distinguish $\Lambda(X)$ from real data pulses, while $\Lambda(X)$ attempts to ‘fool’ $\delta_{T}$. Then $\Lambda(X)$ is fed to $\bar{\Lambda}$ to translate back to the simulated pulse source domain by $\bar{\Lambda}(\Lambda(X))$. A second discriminator $\delta_{S}$ attempts to distinguish $\bar{\Lambda}(\Lambda(X))$ from $X$, while $\bar{\Lambda}$ attempts to ‘fool’ $\delta_{S}$. The $X\rightarrow{}\Lambda(X)\rightarrow{}\bar{\Lambda}(\Lambda(X))$ translation path is termed the forward cycle, denoted by the red arrows in the right-side panel of figure 2. The same process performed in the other direction, starting with the data pulse $X^{\prime}\rightarrow{}\bar{\Lambda}(X^{\prime})\rightarrow{}\Lambda(\bar{\Lambda}(X^{\prime}))$, is called the backward cycle.
3.4 Loss Functions
@@
ensures that the circular translation path $X\rightarrow{}\Lambda(X)\rightarrow{}\bar{\Lambda}(\Lambda(X))$ preserves the original pulse shape. The forward cycle loss shown here corresponds to the case in which a simulated pulse is first fed into REN to produce a data-like pulse; this pulse is then fed through IREN to produce a simulation-like pulse, which should look exactly identical to the input simulated pulse. A similar loss exists for the reciprocal translation path $X^{\prime}\rightarrow{}\bar{\Lambda}(X^{\prime})\rightarrow{}\Lambda(\bar{\Lambda}(X^{\prime}))$.
In each case, the complete cycle should return the original input pulse. The cycle loss calculates the loss associated with this complete cycle. The cycle-consistent loss acts in place of a direct comparison loss to ensure that the underlying physical features that are encoded in a pulse (e.g. multi-site vs. single-site, or energy deposition location) are not destroyed by the translation networks.
@@
A total of eight losses are optimized during training: an adversarial, cycle-consistent, and identity loss for each of the generators, yielding six losses associated with the generators, and an adversarial loss for each discriminator, yielding two losses associated with the discriminators. All losses are computed per pulse and then averaged once per batch. AdamW [29] optimizers are used for all losses. Compared to the traditional Adam optimizer, AdamW applies weight decay as a separate step during gradient descent optimization. This avoids interference with the learning rate schedule and helps stabilize the training process. The six losses for REN and IREN are optimized together, and the discriminators each have their own optimizers. The losses and optimizers are summarized in table 1.
@@
Data Adversarial $L^{\text{adv}}_{\Lambda}$ $-\mathbb{E}_{X}\!\bigl{[}\log(\delta_{T}(\Lambda(X)))\big{]}$
Sim Adversarial $L^{\text{adv}}_{\bar{\Lambda}}$ $-\mathbb{E}_{X^{\prime}}\!\bigl{[}\log(\delta_{S}(\bar{\Lambda}(X^{\prime})))\bigr{]}$
@@
Table 1: Loss terms optimized during CPU-Net training. One optimizer updates both generators using the sum of their six losses, while each discriminator has an independent optimizers.
3.5 Hyperparameter Tuning
@@
baseline_len 200 Samples assigned to the baseline of the pulse.
rising_edge_len 250 Samples assigned to the rising edge of the pulse.
tail_len 350 Samples assigned to the RC decay tail of the pulse.
baseline_weight 3.0 Weight given to the baseline of the pulse.
ris_edge_weight 10.0 Weight given to the rising edge of the pulse.
tail_weight 7.0 Weight given to RC decay tail of the pulse.
@@
lrate_gen $1\times 10^{-3}$ Learning rate for the generator networks.
lrate_disc $1\times 10^{-3}$ Learning rate for the discriminator networks.
cyc_loss_weight 20 Weight of the cycle consistency losses.
iden_loss_weight 5 Weight of the identity loss.
gan_loss_weight 9 Weight of the generator loss.
@@
Table 2: Hyperparameters used for CPU-Net training. These values were obtained through manual tuning and empirical optimization to achieve stable training and good performance.
@@
Weighting the adversarial, cycle-consistency, and identity losses allows us to fine-tune the learning of the generators. At some point, the generator will have learned just enough to fool the discriminator, and further learning might be hampered. The loss weights are used to push the generator to keep improving its translations—maintaining cycle and identity consistency—even after it has been successful in overpowering the discriminator. These hyperparameters were carefully tuned to achieve a balance, as values that are too high will cause the gradient to explode; too low will cause the model not to learn much.
@@
Together, these parameters enable CPU-Net to obtain the right balance during training. A single training takes about 1 GPU hour on an NVIDIA A100 GPU. The trained CPU-Net generates both a REN and an inverse REN, enabling bidirectional translation between the simulation and data domains. The REN is the primary focus of this work.
@@ 4 Results
In order to have accurate simulations, the REN needs to reproduce the correct ensemble distribution of the parameters used in PSD. Since dark matter signatures or $0\nu\beta\beta$ decays would appear as single-site events, experiments often seek to distinguish between single-site and multi-site events using PSD techniques. We trained CPU-Net on $110,000$ FEP pulses. With a batch size of $32$ and training iteration of $7000$, this enables the model to train on every pulse at least twice. We then used $1,200$ SEP events and $3,000$ DEP events for validation. In this section, we present the results of REN’s ability to reproduce detector pulses by comparing the resulting distributions for some critical parameters used in PSD.
@@ 4.1 Training progression
The progression of training losses is shown in figure 4. The cycle-consistent and identity losses converge rapidly towards zero. The adversarial training is evident in the losses for the discriminators and generators. During training, the discriminator initially wins, and its loss decreases while the generator’s loss increases. As training continues, the generator catches up, and the networks approach an equilibrium in which each attempts to out-compete the other. Thus, the loss functions fluctuate as they continuously adapt and improve.
Figure 4: Training losses for CPU-Net. Curves are smoothed using a moving average of 10 samples for clarity. The identity and cycle losses rapidly converged to zero, indicating that the translators learn to preserve the cycle and identity within the first few hundred iterations. The generator and discriminator losses oscillate. This adversarial dynamic helps both models to learn and improve each other.
@@ 4.2 Pulse Translation
Figures 5 and 6 show the SEP pulses as they progress through the cycle in CPU-Net. The network effectively translates the pulses in forward and reverse translations. In the siggen simulations, the RC discharge effect is not modeled, resulting in a flat tail slope of zero. REN learns to transform this flat tail into an exponentially decaying one while including more subtle effects like overshoot and a faster decay early in the tail, matching the observed behavior in real detector data, and IREN learns to transform pulse tails to a zero slope, matching the simulations.
Figure 5: Cycle consistency in pulse translation in the forward direction. Starting with 10 simulated pulses, the REN first translates them into detector-like pulses, and then IREN translates them back to the simulation like pulses.
Figure 6: Cycle consistency in pulse translation in the backward direction. Beginning with 10 detector pulses, the IREN translates them into simulated-like pulses, and then REN translates back to detector-like pulses.
@@ 4.3 Validation of Key pulse Parameters
The accuracy of translations can be assessed at the distribution level by plotting the histogram of key reconstruction parameters. We use the Intersection over Union (IoU) metric, also known as the Jaccard index, to measure the overlap between distributions $A$ and $B$ [30, 31]. We first calculate the histogram of the distribution being compared using the same binning. The intersection is $\sum_{i}\min(\text{bin}_{i}^{A},\text{bin}_{i}^{B})$, the union is $\sum_{i}\max(\text{bin}_{i}^{A},\text{bin}_{i}^{B})$, and
@@ 4.3.1 Drift Time Distribution
Figure 7 illustrates the distribution of $T_{Drift}$ in the SEP and DEP validation datasets. REN successfully learns to slow the drift time of the simulated pulses to match the distribution of the data. The SEP $T_{Drift}$ IoU increases to $62.4\%$ from $39.5\%$, a factor of 1.6 improvement. The DEP IoU increases to $22.5\%$ from $5.4\%$, a factor of 4 improvement.
Figure 7: The distribution of the $T_{Drift}$ on SEP and DEP datasets.
@@ 4.3.2 Current Amplitude
The maximum current amplitude $I_{max}$ is determined by differentiating the pulse and identifying the maximum value of its derivative. For a given event energy, single-site events produce a localized energy deposition, resulting in a sharper and faster increase in $I_{max}$. In contrast, multi-site events, characterized by energy deposition at multiple locations within the detector, yield the same current spread over multiple peaks, decreasing the magnitude of the maximum current value. This distinction makes $I_{max}$ highly effective in differentiating single-site from multi-site events, thereby establishing it as a crucial parameter in pulse shape simulations [13].
Figure 8 illustrates the distribution of $I_{max}$ for both validation datasets. The SEP dataset distributions have two peaks, the higher of which corresponds to single-site events. The peak at a lower value of Imax corresponds to multi-site events where one gamma ray from the positron annihilation, with 511 keV, is fully absorbed at a second site. The DEP distribution has only one peak, corresponding to single-site events. REN learns to correctly slow the current amplitude of the simulated pulses to align with the data, while accurately preserving the difference between single-site and multi-site events. The SEP $I_{max}$ increases to the IoU of $63.71\%$ from $27.53\%$, a factor of 2.3 improvement. The DEP IoU increases to $15.5\%$ from $4.2\%$, a factor of 3.7 improvement over simulated pulses with no applied translation.
Figure 8: Distribution of maximum current amplitude ($I_{max}$) on SEP and DEP validation datasets.

### v2

@@ 3.3 Post Processing
Normalizing and aligning the pulses is crucial for training the network. The simulated pulses are already normalized between 0 and $1$. The raw data pulses are normalized by dividing by the $80\%$ of the average of the last five samples. Pulses are then vertically shifted so that the average of the first $200$ samples is zero. This ensures that all pulses have their preamplifier decay tails and their baselines aligned, which can help the model learn the features better. We used the $99.9\%$ rise time to horizontally align the pulses. We found this method more effective for training the network than aligning by the start of the rising edge, as the pulse shape at the end of the rise time is far more consistent between events than the shape of the beginning of the charge drift, which depends on the position of the energy deposition. The simulated pulses are padded to ensure that there are $400$ samples on both sides of the $99.9\%$ rise time. The preamplifier decay time constant $\tau$ of the data pulses is calculated by the slope of a linear fit of the logarithm of the last $300$ pulse samples. Events with poor fit quality $\chi^{2}$ or anomalous values of $\tau$ are used to identify pile-up pulses in the data and remove them from training.
3.4 Network Training
@@
The training and validation of CPU-Net are conducted in PyTorch [pytorch], following the approach shown schematically in figure 3. We define two PU-based generators called REN and IREN. Then we define two RNN with attention discriminators called $\delta_{S}$ and $\delta_{T}$. Suppose we start with a simulated pulse $X$ and data pulse $X^{\prime}$, with $\Lambda(X)$ the output of the REN generator and $\bar{\Lambda}(X^{\prime})$ the output of the IREN generator. During training, a simulated pulse $X$ is first fed to $\Lambda$ to produce a translated pulse $\Lambda(X)$. The discriminator $\delta_{T}$ attempts to distinguish $\Lambda(X)$ from real data pulses, while $\Lambda(X)$ attempts to ‘fool’ $\delta_{T}$. Then $\Lambda(X)$ is fed to $\bar{\Lambda}$ to translate back to the simulated pulse source domain by $\bar{\Lambda}(\Lambda(X))$. The $X\rightarrow{}\Lambda(X)\rightarrow{}\bar{\Lambda}(\Lambda(X))$ translation path is termed the forward cycle, denoted by the red arrows of figure 3. The same process performed in the other direction, starting with the data pulse $X^{\prime}\rightarrow{}\bar{\Lambda}(X^{\prime})\rightarrow{}\Lambda(\bar{\Lambda}(X^{\prime}))$. A second discriminator $\delta_{S}$ attempts to distinguish $\bar{\Lambda}(X^{\prime})$ from $X$, while $\bar{\Lambda}$ attempts to ‘fool’ $\delta_{S}$. This is called the backward cycle, denoted by the blue arrows of figure 3. The forward and backward cycles are trained simultaneously.
3.4.1 Loss Functions
@@
ensures that the circular translation path $X\rightarrow{}\Lambda(X)\rightarrow{}\bar{\Lambda}(\Lambda(X))$ preserves the original pulse shape. The forward cycle loss shown here corresponds to the case in which a simulated pulse is first fed into REN to produce a data-like pulse; this pulse is then fed through IREN to produce a simulation-like pulse, which should look exactly identical to the input simulated pulse. A similar loss exists for the backward translation path $X^{\prime}\rightarrow{}\bar{\Lambda}(X^{\prime})\rightarrow{}\Lambda(\bar{\Lambda}(X^{\prime}))$.
In each case, the complete cycle should return the original input pulse. The cycle loss calculates the loss associated with this complete cycle. The cycle-consistent loss acts in place of a direct comparison loss to ensure that the underlying physical features that are encoded in a pulse (e.g. multi-site vs. single-site, or energy deposition location) are not destroyed by the translation networks. Since the electronics response effect is not perfectly invertible, this loss will have a natural ‘floor’ associated with the information loss of the true electronics response, even in the case of perfect network performance.
@@
The losses and optimizers are summarized in table 1. A total of eight losses are optimized during training: an adversarial, cycle-consistent, and identity loss for each of the generators, yielding six losses associated with the generators, and an adversarial loss for each discriminator, yielding two losses associated with the discriminators. All losses are computed per pulse and then averaged once per batch. AdamW [adam_w_paper] optimizers are used for all losses. Compared to the traditional Adam optimizer, AdamW applies weight decay as a separate step during gradient descent optimization. This avoids interference with the learning rate schedule and helps stabilize the training process. The six losses for REN and IREN are optimized together in a weighted sum using the weight hyperparameters given in table 2, and the discriminators each have their own optimizers.
@@
Data Adversarial $L^{\text{adv}}_{\Lambda}$ $\mathbb{E}_{X}\!\bigl[\log(\delta_{T}(\Lambda(X)))\big]$
Sim Adversarial $L^{\text{adv}}_{\bar{\Lambda}}$ $\mathbb{E}_{X^{\prime}}\!\bigl[\log(\delta_{S}(\bar{\Lambda}(X^{\prime})))\bigr]$
@@
Table 1: Loss terms optimized during CPU-Net training. One optimizer updates both generators using the weighted sum of their six losses, while each discriminator has an independent optimizer.
3.4.2 Hyperparameter Tuning
@@
baseline_len 200 Samples in the pulse baseline window used for the weighted L1 losses.
rising_edge_len 250 Samples in the pulse rising edge window used for the weighted L1 losses.
tail_len 350 Samples in the pulse decay tail window used for the weighted L1 losses.
baseline_weight 3.0 Weight given to the baseline of the pulse in L1 loss terms.
ris_edge_weight 10.0 Weight given to the rising edge of the pulse in L1 loss terms.
tail_weight 7.0 Weight given to RC decay tail of the pulse in L1 loss terms.
@@
lrate_gen $1\times 10^{-3}$ Initial learning rate for the generator networks.
lrate_disc $1\times 10^{-3}$ Initial learning rate for the discriminator networks.
cyc_loss_weight 20 Weight of the cycle consistency loss terms in the overall generator loss.
iden_loss_weight 5 Weight of the identity loss terms in the overall generator loss.
gan_loss_weight 9 Weight of the adversarial loss terms in the overall generator loss.
@@
Table 2: Hyperparameters used for CPU-Net training. Following an initial hyperparameter space search using Bayesian optimization, these values were obtained through manual tuning to achieve stable training and good performance.
@@
Weighting the adversarial, cycle-consistency, and identity losses allows us to fine-tune the learning of the generators. At some point during the training, the generator will have learned just enough to fool the discriminator, and further learning might be hampered. The loss weights are used to push the generator to keep improving its translations: maintaining cycle and identity consistency—even after it has been successful in overpowering the discriminator. These hyperparameters were carefully tuned to achieve a balance, as values that are too high will cause the gradient to explode; too low will cause the model not to learn much.
@@
Together, these parameters enable CPU-Net to obtain the right balance during training. To optimize the hyperparameters, we conducted an initial search with Bayesian optimization [balandat2020botorch] and then hand-tuned the hyperparameters to reach stable convergence. The hyperparameter optimization was performed with a fixed initial seed. A single training takes about 1 GPU hour on an NVIDIA A100 GPU. The trained CPU-Net generates both a REN and an inverse REN, enabling bidirectional translation between the simulation and data domains. The REN is the primary focus of this work.
@@ 4 Results
Since labeled pairs of simulation and data pulses do not exist, the performance of this approach can only be quantitatively evaluated using information from distribution-level parameters. Individual pulse correspondences are used for a qualitative evaluation, however. We focused our quantitative studies of the results on distributions of pulse shape parameters associated with common analysis techniques used in low-background experiments.
In order to have accurate simulations, the REN needs to reproduce the correct ensemble distribution of the parameters used in PSD. Since dark matter signatures or $0\nu\beta\beta$ decays would appear as single-site events, experiments often seek to distinguish between single-site and multi-site events using PSD techniques. We trained CPU-Net on $110,000$ FEP pulses. With a batch size of $32$ and $7000$ training iterations, this enables the model to train on every pulse at least twice. We then used $12,000$ SEP events and $3,000$ DEP events for validation. In this section, we present the results of REN’s ability to reproduce detector pulses by comparing the resulting distributions for some critical parameters used in PSD.
@@ 4.1 Training Progression
Figure 5 shows the evolution of all eight losses (two for adversarial generators, two for adversarial discriminators, two for cycle consistencies, and two for identities), plotted with a 10-iteration moving average. The cycle and identity losses drop rapidly in the first few hundred iterations and remain small thereafter. The adversarial losses exhibit the expected GAN oscillations: as the discriminators momentarily improve, their losses decrease while the generators’ increase. Then, the trend reverses as the generators adapt. Given that the linear learning-rate decay begins at iteration 1000, and that the discriminators are updated every 30 generator steps, the system enters a regime in which the adversarial losses fluctuate around stable means. The jumps in the losses beyond 3000 iterations, therefore, indicate the alternating advantage typical of adversarial training, not divergence.
Figure 5: Training losses for CPU-Net. Curves are smoothed using a moving average of 10 iterations for clarity. The identity and cycle losses rapidly converge, indicating that the translators learn to preserve the cycle and identity within the first few hundred iterations. Following an initial training phase in which the discriminator dominates, the generator and discriminator losses oscillate. This adversarial dynamic helps both models to learn and improve each other.
@@ 4.2 Pulse Translation
Figures 6(a) and 6(b) show the SEP pulses as they progress through the forward and backward cycles in CPU-Net, respectively. These SEP pulses feature a variety of event topologies, and the network effectively translates the pulses in forward and reverse translations while retaining this information. In the siggen simulations, the RC discharge effect is not modeled, resulting in a flat tail slope of zero. REN learns to transform this flat tail into an exponentially decaying one while including more subtle effects like overshoot and a faster decay early in the tail, matching the observed behavior in real detector data, and IREN learns to transform pulse tails to a zero slope, matching the simulations.
(a) Pulse translation in the forward direction: starting with simulated pulses, the REN first translates them into detector-like pulses, and then IREN translates them back to the simulation-like pulses.
(b) Pulse translation in the backward direction: starting with detector pulses, the IREN translates them into simulation-like pulses, and then REN translates them back to detector-like pulses.
Figure 6: Examples of pulse translation through the CPU-Net cycles for ten single-escape peak pulses.
Figure 7 shows the comparison of one single-site REN-translated pulse with the original siggen simulation input, the siggen simulation with the preamplifier electronics effect enabled, and a comparable data pulse. The two siggen simulated pulses are generated from the same energy deposition information, and the REN-translated pulse is produced from the trained network using the siggen-simulated pulse without the electronics effect enabled as the input. In the absence of labels, there is no way to ensure that the data pulse shown is associated with the same energy deposition location and topology as the simulated pulses, but a single-site data pulse with similar drift time to the simulated pulse has been selected for comparison. All pulses are shown without the pole-zero decay of the tail; the decay is not applied in the two siggen output pulses, and is corrected for by fitting a single decay time constant to the tail in the REN-translated and data pulses. For the siggen simulation with electronics enabled, we applied the built-in siggen preamplifier time constant to the pulse, tuning the time constant to match the mean simulated single-site current amplitude to the mean found in data ($\tau_{rise}=100$ ns).
From figure 7 is is clear that the REN translation provides an improved match relative to the standard simulation approach. Higher-order effects like the pulse overshoot and the smoothing of the initial pulse rise are correctly reproduced by the REN method, but absent from the siggen-produced pulses using the available built-in electronics modeling.
Figure 7: Comparison of a single-site event pulse using siggen with no applied electronics effects, siggen with the preamplifier time constant applied (using $\tau=100$ ns), the REN-translated version of the siggen pulse with no applied electronics effects, and a comparable data pulse. The REN method improves on the match to data, relative to the standard electronics modeling method.
@@ 4.3 Validation of Key Pulse Parameters
The training and evaluation sets are unpaired and there is no known one-to-one correspondence between any detector pulse and any simulated pulse. Therefore, a direct event-level comparison is not possible. We assess agreement statistically at the distribution level for topology–informative parameters: drift time $T_{\mathrm{drift}}$, maximum current amplitude $I_{\max}$, and tail decay constant $\tau$.
We use the Intersection over Union (IoU) metric, also known as the Jaccard index, to measure the overlap between distributions $A$ and $B$ [murphy1996finley, jaccard_index]. We first calculate the histogram of the distribution being compared using the same binning. The intersection is $\sum_{i}\min(\text{bin}_{i}^{A},\text{bin}_{i}^{B})$, the union is $\sum_{i}\max(\text{bin}_{i}^{A},\text{bin}_{i}^{B})$, and
@@ 4.3.1 Drift Time Distribution
Figure 8 illustrates the distribution of $T_{drift}$ in the SEP and DEP validation datasets. REN successfully learns to slow the drift time of the simulated pulses to match the distribution of the data. The SEP $T_{drift}$ IoU increases to $62.3\%$ from $39.5\%$, a factor of 1.6 improvement. The DEP IoU increases to $22.5\%$ from $5.4\%$, a factor of 4 improvement.
Figure 8: The distribution of the $T_{drift}$ on SEP (left) and DEP (right) datasets.
@@ 4.3.2 Current Amplitude
The maximum current amplitude $I_{max}$ is determined by differentiating the pulse and identifying the maximum value of its derivative. For a given event energy, single-site events produce a localized energy deposition, resulting in a sharper and faster increase in $I_{max}$. In contrast, multi-site events, characterized by energy deposition at multiple locations within the detector, yield the same current spread over multiple peaks, decreasing the magnitude of the maximum current value. This distinction makes $I_{max}$ highly effective in differentiating single-site from multi-site events, thereby establishing it as a crucial parameter in pulse shape simulations [AvsE].
Figure 9 illustrates the distribution of $I_{max}$ for both validation datasets. The SEP dataset distributions have two peaks, the higher of which corresponds to single-site events. The peak at a lower value of Imax corresponds to multi-site events where one gamma ray from the positron annihilation, with 511 keV, is fully absorbed at a second site. The DEP distribution has only one peak, corresponding to single-site events. REN learns to correctly slow the current amplitude of the simulated pulses to align with the data, while accurately preserving the difference between single-site and multi-site events. The SEP $I_{max}$ increases to the IoU of $63.7\%$ from $27.5\%$, a factor of 2.3 improvement. The DEP IoU increases to $15.5\%$ from $4.2\%$, a factor of 3.7 improvement over simulated pulses with no applied translation.
Figure 9: Distribution of maximum current amplitude ($I_{max}$) on SEP (left) and DEP (right) validation datasets.